OpenStudy (anonymous):

Prime or Composite 153 and 201 and 57

OpenStudy (agreene):

All of these are Composite. Like I mentioned previously, all even numbers are divisible by 2. The majority of odds are divisible by 3. So, when checking for Primes I usually just start by seeing if its odd. If it is I divide by 3, if that isn't evenly divisible (no multiple of 3) then I try some harder methods. As it happens all 3 of these are divisible by 3. 153/3 = 51 so 51*3 = 153 Composite 201/3 = 67 so 67*3 = 201 Composite 57/3 = 19 so 19*3 = 57 Composite